# Strong Coupling of Hydrodynamics and Reactions in Nuclear Statistical Equilibrium for Modeling Convection in Massive Stars

### v1

@@ 5.2.2 Convective Initialization
We map the initial model onto the Cartesian grid by taking pressure, density, composition, and $Y_{e}$ as inputs and interpolating them to cell centers. We then check if a cell is in NSE, and if so, we update the composition and $\langle B/A\rangle$ from the NSE table; otherwise, we compute $\langle B/A\rangle$ using the composition in the cell. Finally, we call the EOS to get the temperature. Just like with the initial model, if we are in NSE, then we need to do the EOS inversion self-consistent with the NSE table, analogous to Eq. 44, but using the interpolated $p$ instead of $e$.
@@ 5.2.3 Simulations
ms-enu $8192\times 16384$ $1024\times 2048$ 2: $\times,2\times$ 2 × 0.4 50 $10^{3}$ includes $\epsilon_{\nu,\mathrm{weak}}$
@@
We run 7 simulations, each with slightly different parameters to assess the sensitivity of the evolution to the problem setup. Table  3 summarizes the simulations. The default simulation, ms, uses a domain that is $8.192\times 10^{9}~{}\mathrm{cm}\times 1.6384\times 10^{9}~{}\mathrm{cm}$ in a 2D axisymmetric ($r$-$z$) geometry. The base grid is $1024\times 2048$ zones and 2 levels of refinement are used, each jumping by a factor of 2, giving a fine-grid resolution of 20 km. This is slightly less than the 16 km resolution of Couch et al. (2015). Refinement is done by tagging all cells that have a density larger than $10^{4}~{}\mathrm{g~{}cm^{-3}}$, encompassing most of the O shell. Figure 1 shows the extent of the refined region and the domain. Subcycling is used to advance coarse zones at larger timesteps than fine zones. The gravitational acceleration is computed as a simple monopole, with the source term treated conservatively (as described in Katz et al. 2016). Integration is done with CFL = 0.4, without any reaction-based timestep limiter. Evolving to $t-\tau_{\mathrm{drive}}=300s$ takes 98755 coarse timesteps for ms.
@@
The contribution to the energy generation from weak-rate neutrinos, $\epsilon_{\nu,\mathrm{weak}}$, is disabled—as we’ll see shortly, including this term leads to an almost instantaneous collapse, suggesting a mismatch between the weak rate neutrino physics in the MESA model and what we’ve modeled here. Consistency between the nuclear physics used in the 1D model and that used here will be explored in a follow-on paper. The convection is driven for $\tau_{\mathrm{drive}}=50~{}\mathrm{s}$, after which the problem is left to evolve.
@@
Figure 2: Mach number sequence for the ms simulation focusing on the region near the core. The first frame is just after the drive initial convection phase.
Figure 3: Time sequence of $\bar{A}$ for the ms simulation, focusing on the region near the core.
@@
Figure 4: Time sequence of the energy generation rate for the ms simulation focusing on the region near the core. The back contour separates the region where the NSE table is used from the region where the full network is integrated.
@@
Zooming out, Figure 5 shows the state on a much larger scale. Here we clearly see the large scale vortices in the O shell that dominate the flow. We also see that the Mach number can get quite high there. Again, we expect much of this behavior to be an artifact of the 2D geometry.
@@
Figure 5: The state of the convection in the ms simulation in a large region of the star, 200 s after the drive initial convection. The full O shell is seen on this scale.
@@
Next we look at how the integral properties of the simulations differ as we adjust some model parameters. The ms-100 simulation doubles $\tau_{\mathrm{drive}}$ from $50~{}\mathrm{s}$ to $100~{}\mathrm{s}$. The ms-cfl2 and ms-cfl8 simulations explore CFL numbers of 0.2 and 0.8, respectively. The ms-sp simulation uses a domain that is 50% larger (in each dimension) and drops $\rho_{\mathrm{sponge}}$ from $10^{3}~{}\mathrm{g~{}cm^{-3}}$ to $100~{}\mathrm{g~{}cm^{-3}}$—this allows us to understand how well our outer boundary works. This also means that the evolution of the full C shell is now captured on the grid. Figure 6 shows the evolution of the maximum density on the grid as a function of time for the different simulations. We see strong agreement between ms, ms-100, ms-cfl2, ms-cfl8, and ms-sp—for each, the central density slowly increases in time, reaches a peak around 150 s, and then begins to decrease. This peak corresponds to the point shown in Figure 3 where the O convection begins to erode the Si layer. Again we note that this is very likely a 2D effect. Figure 7 shows the peak temperature evolution, and we see the same trend for this group of simulations. Figure 8 shows the minimum $Y_{e}$ on the grid. All simulations show a decrease in time, and none fall below $Y_{e}=0.43$, which is the boundary of our table. Finally, Figure 9 shows the total mass of ${}^{28}\mathrm{Si}$ and ${}^{32}\mathrm{S}$. We see the mass of these elements decreasing initially, but again, near 150 s, the trend inverts and we begin producing more of these nuclei. This is consistent with the idea that the strong 2D O-shell convection begins disrupting the Si shell, bringing fresh O into hot regions and increases these abundances. The key takeaway from these figures is that our evolution is largely insensitive to the domain size / sponging, CFL, and initialization.
@@
Finally the ms-enu simulation enables the $\epsilon_{\nu,\mathrm{weak}}$ term in the energy generation in the NSE region. We see that the density increases dramatically right from the start, leading to a collapse. This suggests that the MESA initial model, which used aprox21, did not capture all of these neutrino losses.
@@
Figure 6: Peak density vs. time since the drive initial convection process ended for the 7 different simulations.
Figure 7: Peak temperature vs. time since the drive initial convection process ended for the 7 different simulations.
Figure 8: Minimum $Y_{e}$ vs. time since the drive initial convection process ended for the 7 different simulations.
Figure 9: Total ${}^{28}\mathrm{Si}$ and ${}^{32}\mathrm{S}$ mass vs. time since the drive initial convection process ended for the 7 different simulations.
@@ 6 Summary
We presented a simplified spectral deferred corrections scheme for coupling hydrodynamics and reactions when nuclear statistical equilibrium is established. We showed that the method is second-order accurate and provides an effective simulation method for modeling convection in massive stars, including the evolution of the core on the grid. Furthermore, the time-integration algorithm allows for evolution on the hydrodynamics timescale and with GPU acceleration is very efficient on today’s supercomputers. This entire simulation methodology is freely available in the AMReX-Astro github organization [^1] https://github.com/AMReX-Astro/.
@@
We will follow this work with a 3D simulation. Currently we are fully refining all regions where the density is greater than $10^{4}~{}\mathrm{g~{}cm^{-3}}$. This corresponds to a radius of about 20,000 km. In Couch et al. (2015), only the inner 2,500 km was fully refined. To reduce the computational cost for 3D, we will need to relax our refinement with radius compared to the 2D models presented here. Initial estimates show that a 3-level simulation, evolving $\rho>3\times 10^{5}~{}\mathrm{g~{}cm^{-3}}$ at 20 km (corresponding to $r<4700~{}\mathrm{km}$), $\rho>2\times 10^{3}~{}\mathrm{g~{}cm^{-3}}$ at 80 km, and the base grid at 320 km resolution (each jump is 4×4$\times$4 ×) will need approximately 40,000 node hours on the OLCF Frontier machine to evolve for 350 s (including the 50 s of convective initialization). This estimate uses HIP to offload the compute kernels to the AMD Instinct MI250X GPUs, similar to how we use CUDA with NVIDIA GPUs at NERSC. We expect to bring this cost down with further optimization on AMD GPUs.
@@
The focus of this paper was the development of an accurate time-integration strategy for convective burning in massive stars. In follow-on work, we will explore the physics of this convective burning and the evolution of different progenitor massive stars. We will also explore how sensitive the numerical results are to our choice of network (both the on-grid network and our NSE tabulations), spatial resolution, and the details of the initial model. For this new work, we will generate new initial models that start earlier in evolution and explore getting better consistency with the nuclear physics inputs in the stellar evolution code.
@@ Appendix A Constructing the NSE Table
We use pynucastro (Willcox & Zingale, 2018; Smith Clark et al., 2022) to construct the NSE table, making use of the features in the latest version (Smith Clark et al., 2023). We use the weak rates from Langanke & Martínez-Pinedo (2001) to be consistent with the work from Seitenzahl et al. (2009); Ma et al. (2013). But by using a completely open, reproducible framework built from pynucastro, we can regenerate the table easily as newer rate compilations are added to pynucastro, for example the those of Giraud et al. (2022). The solution procedure follows that of Seitenzahl et al. (2009), and we note that we assume that the nuclei behave as an ideal gas, consistent with our equation of state and a good approximation for the densities we model here.
Nuclear masses are taken the atomic mass evaluation (AME) 2016 database (Huang et al., 2017; Wang et al., 2017) and spins from Kondev et al. (2021). The temperature-dependent partition functions come from Rauscher et al. (1997); Rauscher (2003). Finally, we use the Coulomb screening implementation from Chabrier & Potekhin (1998).
@@
$\epsilon_{\nu}=\sum_{r\in\mathrm{rates}}Y_{r}{\epsilon_{\nu}}_{r}(\rho Y_{e},T)$ (A9)

### v2

@@ 5.2.2 Convective Initialization
We map the initial model onto the Cartesian grid by taking pressure, density, composition, and $Y_{e}$ as inputs and interpolating them to cell centers. We then check if a cell is in NSE, and if so, we update the composition and $\langle B/A\rangle$ from the NSE table; otherwise, we compute $\langle B/A\rangle$ using the composition in the cell. Finally, we call the EOS to get the temperature. Just like with the initial model, if we are in NSE, then we need to do the EOS inversion self-consistent with the NSE table, analogous to Eq. 44, but using the interpolated $p$ instead of $e$. For our domain size, the mass on the grid is $2.7~{}M_{\odot}$.
@@ 5.2.3 Simulations
ms-enu $8192\times 16384$ $1024\times 2048$ 2: $\times,2\times$ 2 × 0.4 50 $10^{3}$ includes $\epsilon_{\nu,\mathrm{react}}$
@@
We run 7 simulations, each with slightly different parameters to assess the sensitivity of the evolution to the problem setup. Table  3 summarizes the simulations. The default simulation, ms, uses a domain that is $8.192\times 10^{9}~{}\mathrm{cm}\times 1.6384\times 10^{9}~{}\mathrm{cm}$ in a 2D axisymmetric ($r$-$z$) geometry. The base grid is $1024\times 2048$ zones and 2 levels of refinement are used, each jumping by a factor of 2, giving a fine-grid resolution of 20 km. This is slightly coarser than the 16 km resolution of Couch et al. (2015). Refinement is done by tagging all cells that have a density larger than $10^{4}~{}\mathrm{g~{}cm^{-3}}$, encompassing most of the O shell. Figure 1 shows the extent of the refined region and the domain. Subcycling is used to advance coarse zones at larger timesteps than fine zones. The gravitational acceleration is computed as a simple monopole, with the source term treated conservatively (as described in Katz et al. 2016). Integration is done with CFL = 0.4, without any reaction-based timestep limiter. Evolving to $t-\tau_{\mathrm{drive}}=300s$ takes 98755 coarse timesteps for ms.
@@
The contribution to the energy generation from weak-rate neutrinos, $\epsilon_{\nu,\mathrm{react}}$, is disabled—as we’ll see shortly, including this term leads to an almost instantaneous collapse, suggesting a mismatch between the weak rate neutrino physics in the MESA model and what we’ve modeled here. The initial model we are using was run in MESA with the aprox21 network, which does not include all of these neutrinos. Consistency between the nuclear physics used in the 1D model and that used here will be explored in a follow-on paper. The convection is driven for $\tau_{\mathrm{drive}}=50~{}\mathrm{s}$, after which the problem is left to evolve.
@@
Figure 2: Mach number sequence for the ms simulation focusing on the region near the core. The first frame is just after the drive initial convection phase. The black contour separates the region where the NSE table is used from the region where the full network is integrated.
@@
Figure 3: Time sequence of $\bar{A}$ for the ms simulation, focusing on the region near the core. The black contour separates the region where the NSE table is used from the region where the full network is integrated.
Figure 4: Time sequence of the energy generation rate for the ms simulation focusing on the region near the core. The black contour separates the region where the NSE table is used from the region where the full network is integrated.
@@
Zooming out, Figure 5 shows the state on a much larger scale (four times larger than Figures 24). A gray contour is added that marks $X({}^{16}\mathrm{O})=0.5$, which both demonstrates the extent of the oxygen shell and that there is mixing near the core resulting in pockets depleted of oxygen. We also clearly see the large scale vortices in the O shell that dominate the flow, and note that the Mach number can get quite high there. Again, we expect much of this behavior to be an artifact of the 2D geometry.
@@
Figure 5: The state of the convection in the ms simulation in a large region of the star, 200 s after the drive initial convection. The full O shell is seen on this scale. The gray contour marks an ${}^{16}\mathrm{O}$ mass fraction of $0.5$.
@@
Next we look at how the integral properties of the simulations differ as we adjust some model parameters. The ms-100 simulation doubles $\tau_{\mathrm{drive}}$ from $50~{}\mathrm{s}$ to $100~{}\mathrm{s}$. The ms-cfl2 and ms-cfl8 simulations explore CFL numbers of 0.2 and 0.8, respectively. The ms-sp simulation uses a domain that is 50% larger (in each dimension) and drops $\rho_{\mathrm{sponge}}$ from $10^{3}~{}\mathrm{g~{}cm^{-3}}$ to $100~{}\mathrm{g~{}cm^{-3}}$—this allows us to understand how well our outer boundary works. This also means that the evolution of the full C shell is now captured on the grid. Figure 6 shows the evolution of the maximum density on the grid as a function of time for the different simulations as well as the difference with respect to the ms simulation for ms-100, ms-cfl2, ms-cfl8, and ms-sp. For those four simulations, the central density slowly increases in time, reaches a peak around 150 s, and then begins to decrease. This peak corresponds to the point shown in Figure 3 where the O convection begins to erode the Si layer. Again we note that this is very likely a 2D effect. The largest relative difference in density is less than 1% for these simulations, demonstrating strong agreement as we vary the simulation parameters. Figure 7 shows the peak temperature evolution and difference with respect to ms, and we see the same trend for this group of simulations,with the maximum difference around 0.5%. Figure 8 shows the minimum $Y_{e}$ on the grid. All simulations show a decrease in time, and none fall below $Y_{e}=0.43$, which is the boundary of our table. The comparison with ms shows that the largest difference in the global $Y_{e}$ minimum is less than 1%. Finally, Figure 9 shows the total mass of ${}^{28}\mathrm{Si}$ and ${}^{32}\mathrm{S}$. We see the mass of these elements decreasing initially, but again, near 150 s, the trend inverts and we begin producing more of these nuclei. This is consistent with the idea that the strong 2D O-shell convection begins disrupting the Si shell, bringing fresh O into hot regions and increases these abundances. Once again, the comparison with the ms simulation shows strong agreement across the variations, with the largest difference around 1%. The key takeaway from these figures is that our evolution is largely insensitive to the domain size / sponging, CFL, and initialization.
@@
Finally the ms-enu simulation enables the $\epsilon_{\nu,\mathrm{react}}$ term in the energy generation in the NSE region. We see that the density increases dramatically right from the start, leading to a collapse. This suggests that the MESA initial model, which used aprox21, did not capture all of these neutrino losses.
@@
Refer to captionRefer to caption
Figure 6: (left) Peak density vs. time since the drive initial convection process ended for the 7 different simulations. (right) Relative difference in peak density with the ms simulation for several cases.
Refer to captionRefer to caption
Figure 7: (left) Peak temperature vs. time since the drive initial convection process ended for the 7 different simulations. (right) Relative difference in peak temperature with the ms simulation for several cases.
Refer to captionRefer to caption
Figure 8: (left) Minimum $Y_{e}$ vs. time since the drive initial convection process ended for the 7 different simulations. (right) Relative difference in minimum $Y_{e}$ with the ms simulation for several cases.
Refer to captionRefer to caption
Figure 9: Sum of the ${}^{28}\mathrm{Si}$ and ${}^{32}\mathrm{S}$ mass vs. time since the drive initial convection process ended for the 7 different simulations. (right) Relative difference in the total ${}^{28}\mathrm{Si}$ and ${}^{32}\mathrm{S}$ mass with the ms simulation for several cases.
@@ 6 Summary
The focus of this paper was the development of an accurate time-integration strategy for reactive flows that enter nuclear statistical equilibrium. We showed that our new simplified spectral deferred corrections scheme for coupling hydrodynamics, reactions, and NSE is second-order accurate and provides an effective simulation method for modeling convection in massive stars, including the evolution of the core on the grid. Furthermore, the time-integration algorithm allows for evolution on the hydrodynamics timescale, and our tests investigating the timestep size and initialization procedure showed that the simulations are robust, with only small (< 10%) variations. Furthermore, with GPU acceleration, the new algorithm is very efficient on today’s supercomputers. This entire simulation methodology is freely available in the AMReX-Astro github organization [^1] https://github.com/AMReX-Astro/.
@@
This work is the basis for a study of convection in massive stars in 3D. Currently we are fully refining all regions where the density is greater than $10^{4}~{}\mathrm{g~{}cm^{-3}}$. This corresponds to a radius of about 20,000 km. In Couch et al. (2015), only the inner 2,500 km was fully refined. To reduce the computational cost for 3D, we will need to relax our refinement with radius compared to the 2D models presented here. Initial estimates show that a 3-level simulation, evolving $\rho>3\times 10^{5}~{}\mathrm{g~{}cm^{-3}}$ at 20 km (corresponding to $r<4700~{}\mathrm{km}$), $\rho>2\times 10^{3}~{}\mathrm{g~{}cm^{-3}}$ at 80 km, and the base grid at 320 km resolution (each jump is 4×4$\times$4 ×) will need approximately 40,000 node hours on the OLCF Frontier machine to evolve for 350 s (including the 50 s of convective initialization). This estimate uses HIP to offload the compute kernels to the AMD Instinct MI250X GPUs, similar to how we use CUDA with NVIDIA GPUs at NERSC. We expect to bring this cost down with further optimization on AMD GPUs.
@@ Appendix A Constructing the NSE Table
We use pynucastro (Willcox & Zingale, 2018; Smith Clark et al., 2022) to construct the NSE table, making use of the features in the latest version (Smith Clark et al., 2023). We use the weak rates from Langanke & Martínez-Pinedo (2001) to be consistent with the work from Seitenzahl et al. (2009); Ma et al. (2013). By using a completely open, reproducible framework built from pynucastro, we can regenerate the table easily as newer rate compilations are added to pynucastro, for example the those of Giraud et al. (2022). The solution procedure follows that of Seitenzahl et al. (2009), and we note that we assume that the nuclei behave as an ideal gas, consistent with our equation of state and a good approximation for the densities we model here.
Nuclear masses are taken from the atomic mass evaluation (AME) 2016 database (Huang et al., 2017; Wang et al., 2017) and spins from Kondev et al. (2021). The temperature-dependent partition functions come from Rauscher et al. (1997); Rauscher (2003). Finally, we use the Coulomb screening implementation from Chabrier & Potekhin (1998).
@@
$\epsilon_{\nu,\mathrm{react}}=\sum_{r\in\mathrm{rates}}Y_{r}{\epsilon_{\nu}}_{r}(\rho Y_{e},T)$ (A9)